can you help me?
1.an older computer unit requires 12min more to print a company's payroll than does a newer computer unit. used together the computers can print the payroll in 8 min. find the time the new computer unit, used alone, will require to print the payroll.
2.working together. Dan and Leah can mow a lawn in 1 h and 12 min. Working alone, Leah can mow the lawn in 1 h less time than it would take Dan. Find the time it takes for each of them, working on his/her own, to mow the lawn.
i hope you can help me..
god bless..

Answer by scott8148(6628)

when people (or machines) combine to do a task; each does some fraction of the task based on their individual speed

(combined speed)/(individual speed)=individual fraction of task


1. let x="new time" ___ (8/12)+(8/x)=1 ___ multiplying by 12x gives 8x+96=12x

2. working in minutes ___ let x=Leah, so x+60=Dan ___ (72/x)+(72/(x+60))=1
___ multiplying by (x)(x+60) gives 72(x+60)+72(x)=(x)(x+60) ___ 144x+4320=x^2+60x
___ 0=x^2-84x-4320 ___ 0=(x-120)(x+36)